# Vertex ordering characterizations of interval $r$- graphs

### v1

@@ Abstract
An $r$-partite graph is an interval $r$-graph if corresponding to each vertex we can assign an interval of the real line such that two vertices $u$ and $v$ of different partite sets are adjacent if and only if their corresponding intervals intersect. In this paper, we provide two vertex-ordering characterizations of interval $r$-graphs and identify forbidden patterns for interval $r$-graphs in terms of specific orderings of their vertices.
@@ 1 Introduction
A graph $G=(V,E)$ is called an interval graph, if there exists a family $\mathcal{I}=\{I_{v}:v\in V\}$, of intervals of the real line such that, for $a,b\in V$ the vertices $a$ and $b$ are adjacent in $G$ if and only if $I_{a}\cap I_{b}\neq\emptyset$. A bigraph $B=(X,Y,E)$ is said to be an interval bigraph if there exists a family $\mathcal{I}=\{I_{v}:v\in X\cup Y\}$ of intervals of the real line such that for all $x\in X$ and $y\in Y$ the vertices $x$ and $y$ are adjacent in $B$ if and only if $I_{x}\cap I_{y}\neq\emptyset$. Interval bigraphs were introduced in [5] and have been studied in [2, 3, 6, 7]. Interval digraphs were introduced by Sen et al. [9]. It was shown in [3] that the notions of interval digraphs and interval bigraphs are equivalent. Interval bigraphs or interval digraphs have become of interest in such new areas as graph homomorphism, e.g [4]. The concept of interval bigraph was generalized by Brown [1]. A graph $G=(V,E)$ is called interval $r$-graph if the vertex set $V$ can be partitioned into $r$ stable sets say $V_{1}$, $V_{2}$,…, $V_{r}$ and corresponding to each vertex we can assign an interval such that two vertices $u$ and $v$ are adjacent in $G$ if and only if the corresponding intervals $I_{u}$ and $I_{v}$ have non empty intersection and $u$ and $v$ belong to different partite set of vertices.
@@ 2 Main Result
Case 1: $i>j$.
From the interval representation (see Figure 2), any vertex appearing between $v_{j}$ and $v_{i}$ in this ordering, and belonging to a partite set other than $X_{\beta}$, will be adjacent to $v_{j}$. Consequently, $C_{j}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
@@
Case 2: $i<j$.
Similarly, any vertex appearing between $v_{i}$ and $v_{j}$ that does not belong to the same partite set $X_{\alpha}$ will be adjacent to $v_{i}$. Hence, $R_{i}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
@@
Example 1. Consider the following $3$-partite graph. Using the above theorem we show that it is an interval $3$-graph.
@@
Figure 5: The adjacency matrix of the 33 partite graph in Figure 2, where the rows and columns are arranged according to the increasing order of their indices, and corresponding $\mathcal{R}_{i}$’s and $\mathcal{C}_{j}$’s are shown in the elliptic regions. And its interval representation is as follows:
Red: $I_{v_{1}}=[1,3]$, $I_{v_{6}}=[6,9]$, $I_{v_{10}}=[10,10]$
Green: $I_{v_{2}}=[3,4]$, $I_{v_{8}}=[8,10]$, $I_{v_{9}}=[9,9]$
Black: $I_{v_{3}}=[3,3]$, $I_{v_{4}}=[4,6]$, $I_{v_{5}}=[5,8]$, $I_{v_{7}}=[7,9]$.
@@
Figure 8: Forbidden patterns.
@@
  • Case 1. $v_{j}$ and $v_{k}$ belong to the same partite set $X_{\gamma}$. Then configuration (i) (Figure 8) occurs — a contradiction.
@@
    • Subcase 1: If $v_{j}$ is not adjacent to $v_{k}$, then configuration (ii) occurs — a contradiction.
    • Subcase 2: If $v_{j}$ is adjacent to $v_{k}$, then configuration (iii) occurs — a contradiction.
@@ Acknowledgement
The first author sincerely acknowledges the Council of Scientific and Industrial Research (CSIR), India, for providing financial support through the CSIR Fellowship (File No. 09/0028
(11986)/2021-EMR-I).

### v2

@@ Abstract
An $r$-partite graph is an interval $r$-graph if corresponding to each vertex we can assign an interval of the real line such that two vertices $u$ and $v$ of different partite sets are adjacent if and only if their corresponding intervals intersect. In this paper, we provide two vertex-ordering characterizations of interval $r$-graphs, utilizing generalized interval orderings and $r$-interval orderings, and identify forbidden patterns for interval $r$-graphs in terms of specific orderings of their vertices.
@@ 1 Introduction
A graph $G=(V,E)$ is called an interval graph, if there exists a family $\mathcal{I}=\{I_{v}:v\in V\}$, of intervals of the real line such that, for $a,b\in V$ the vertices $a$ and $b$ are adjacent in $G$ if and only if $I_{a}\cap I_{b}\neq\emptyset$. A bigraph $B=(X,Y,E)$ is said to be an interval bigraph if there exists a family $\mathcal{I}=\{I_{v}:v\in X\cup Y\}$ of intervals of the real line such that for all $x\in X$ and $y\in Y$ the vertices $x$ and $y$ are adjacent in $B$ if and only if $I_{x}\cap I_{y}\neq\emptyset$. Interval bigraphs were introduced in [5] and have been studied in [2, 3, 6, 7]. Interval digraphs were introduced by Sen et al. [8]. It was shown in [3] that the notions of interval digraphs and interval bigraphs are equivalent. Interval bigraphs or interval digraphs have become of interest in such new areas as graph homomorphism, e.g [4]. The concept of interval bigraph was generalized by Brown [1]. A graph $G=(V,E)$ is called interval $r$-graph if the vertex set $V$ can be partitioned into $r$ stable sets say $V_{1}$, $V_{2}$,…, $V_{r}$ and corresponding to each vertex we can assign an interval such that two vertices $u$ and $v$ are adjacent in $G$ if and only if the corresponding intervals $I_{u}$ and $I_{v}$ have non-empty intersection and $u$ and $v$ belong to different partite sets of vertices.
@@ 2 Main Result
Case 1:
When $i>j$. From the interval representation (see Figure 2), any vertex appearing between $v_{j}$ and $v_{i}$ in this ordering, and belonging to a partite set other than $X_{\beta}$, will be adjacent to $v_{j}$. Consequently, $C_{j}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
@@
Case 2:
When $i<j$. Similarly, any vertex appearing between $v_{i}$ and $v_{j}$ that does not belong to the same partite set $X_{\alpha}$ will be adjacent to $v_{i}$ (see Figure 3). Hence, $R_{i}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
@@
Example 1. Consider the following $3$-partite graph in Figure 4. Using the above theorem we show that it is an interval $3$-graph.
@@
Figure 5: The adjacency matrix of the 3-partite graph in Figure 4, where the rows and columns are arranged according to the increasing order of their indices, and corresponding $\mathcal{R}_{i}$’s and $\mathcal{C}_{j}$’s are shown in the elliptic regions. Its interval representation is as follows: Red: $I_{v_{1}}=[1,3]$, $I_{v_{6}}=[6,9]$, $I_{v_{10}}=[10,10]$; Green: $I_{v_{2}}=[3,4]$, $I_{v_{8}}=[8,10]$, $I_{v_{9}}=[9,9]$; Black: $I_{v_{3}}=[3,3]$, $I_{v_{4}}=[4,6]$, $I_{v_{5}}=[5,8]$, $I_{v_{7}}=[7,9]$.
@@
Figure 8: Forbidden patterns for interval $r$-graphs ($r\geq 3$) in terms of specific orderings of their vertices.
@@
  • Case 1. $v_{j}$ and $v_{k}$ belong to the same partite set $X_{\gamma}$. Then configuration (i) (see Figure 8) occurs — a contradiction.
@@
    • Subcase 1: If $v_{j}$ is not adjacent to $v_{k}$, then configuration (ii) (see Figure 8) occurs — a contradiction.
    • Subcase 2: If $v_{j}$ is adjacent to $v_{k}$, then configuration (iii) (see Figure 8) occurs — a contradiction.
@@ Acknowledgement
The first author sincerely acknowledges the Council of Scientific and Industrial Research (CSIR), India [File No. 09/0028(11986)/2021-EMR-I], for providing financial support through a fellowship to pursue the Ph.D. degree.